[figure: 9.7 cm]

MAGNETIZATION REVERSAL PROCESSES IN ULTRATHIN

Co/Pt FILMS

S. Lemerle^{a}, M. Kisielewski^{b}, A. Maziewski^{b}, C. Chappert^{c}, V. Mathet^{c}, and J. Ferre^{a}

aLaboratoire de Physique des Solides, Universite Paris-Sud, 91405 Orsay, France bInstitute of Physics, University at Bialystok, 15-424 Bialystok, Lipowa 41, Poland cInstitut d’Electronique Fondamentale, Universite Paris-Sud, 91405 Orsay, France

Magnetization reversal processes in ferromagnetic ultrathin Pt/Co/Pt structures, deposited on Si/SiN substrates, were investigated by polar magneto-optical Kerr magnetometry. The spin reorientation transition takes place for a Co thickness t=1.6nm; for t<1.6nm the easy anisotropy axis lies perpendicular to the film while it falls in-plane at larger thickness. We have limited the present study to samples with t<1.4nm, which exhibit square histeresis loops. The in-field magnetization processes and their dynamics, starting from saturated or demagnetized state, obtained under a perpendicular magnetic field, have been studied for different field directions. In particular, the magnetic aftereffect strongly depends on the field configuration. Numerical simulations, similar to these developed in Ref.[1], allow to describe the magnetic non-reversibility as related to time dependent effects. Experimental data have been succesfully compared with those obtained previously [1] on Au/Co/Au films. The in-plane field magnetization processes investigation was used for the magnetic anisotropy study. Work partially supported by ESF NANOMAG Program & Polish Grant No 2P03B06515. J. Ferre et al., Phys.Rev. B 55, 15092 (1997)

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : M.Kisielewski

Address for correspondence : Institute of Physics, University at Bialystok 15-424 Bialystok, Lipowa 41 Poland

Email address : kisiel@cksr.ac.bialystok.pl